# Dynamical patterns and nonreciprocal effective interactions in an active-passive mixture through exact hydrodynamic analysis

### v2

@@ 3 Traveling solutions in large systems
$-(c/L)\varrho+F&=A_{1},&\qquad F&=-\varrho^{\prime}-\tfrac{\text{Pe}^{2}}{2}(1-\varrho)\partial_{z}[d_{s}(\varrho)\varrho_{a}]\\ -(c/L)\varrho_{a}+F_{a}&=A_{2},&F_{a}&=-d_{s}(\varrho)\varrho_{a}^{\prime}-\varrho_{a}{\cal D}(\varrho)\varrho^{\prime}-\tfrac{\text{Pe}^{2}}{2}\left[\varrho_{a}s(\varrho)+d_{s}(\varrho)\right]\partial_{z}[d_{s}(\varrho)\varrho_{a}],\end{aligned}$ (S68)
@@
[…]
@@
This yields a set of $32N+$ 3 unknowns ($\varrho_{i},\varrho_{a,i},c,A_{1},A_{2}$) and $22N+$ 2 equations [Eqs. (S70) and (S71)]. As in Subsec. 2, the additional constraint comes from noting that the problem has translational symmetry; we fix $\varrho_{0}=\phi$.
@@
Figure S5: CP solution to (1). The initial condition is the homogeneous state ($\rho_{\pm}=\phi_{a}/2,\rho_{0}=\phi_{p}$) perturbed by a uniform random perturbation (see Sec. 1 for details). (a) Density profile at $t=4000$. (b) Kymographs showing the spatiotemporal dynamics. Parameters: $\text{Pe}=7.5$, $L=100.0$, $\phi_{a}=0.5$, $\phi_{p}=0.3$, $\Delta x=0.05$.
@@
Figure S6: Two solutions, both obtained from Eq. (8) with the same parameters. (a) T solution. (b) PS solution. The top panels display the solution profile as a function of $z/L$. The bottom panels display the profile (solid black) overlayed on phase diagrams spanned by $\phi_{a}$ and $\phi_{p}$. The Spinodal (Blue) encloses the region of linear stability of homogeneous solutions, and the Binodal (Red) encloses the region of Phase separation. Black crosses mark the volume fraction. Parameters: $\text{Pe}=7.5$, $\phi_{a}=0.4608$, $\phi_{p}=0.2201$, $N=1024$
@@
Figure S7: (a) Sketch of T solution with positive velocity, showing density of active and passive particles as a function of position (see Fig. 4 of main text for a similar solution with negative velocity). The red arrows indicate the self-propulsive force acting on the active particles. The blue arrows indicate forces applied by the active particles on the passive ones (via the cross-diffusion terms in the equation of motion). This generates the positive propagation. (b) Sketch of a CP solution at three different times in a large system (see Fig. 3 of main text for a similar solution in a smaller system). Initially the CP solution is well-approximated by a combination of two T solutions with opposite velocities of equal magnitude. The two propagating clusters approach each other and eventually collide. After the collision process one again sees a combination of two T solutions which travel away from each other. The cycle will eventually repeat due to periodic boundaries.
@@
For the outer1 problem with parameters $\phi$ and $\nu$, we first solve the coexisting phases problem (Sec. C) given $\nu$. This results in vapor and liquid phases $\phi_{v}(\nu)$ and $\phi_{l}(\nu)$, respectively, as well as individual components $(\varrho_{\sigma})_{v}$ and $(\varrho_{\sigma})_{l}$ in each phase. They satisfy $(\varrho_{\sigma})_{v}\equiv(\varrho_{+},\varrho_{-},\varrho_{0})_{v}=(\tfrac{1}{2}\phi_{a,g},\tfrac{1}{2}\phi_{a,g},\phi_{p,g})$ and similarly for $(\varrho_{\sigma})_{l}$. If $\phi_{v}(\nu)=\phi_{l}(\nu)$, it means there is no inner region and outer1 has no solution (and we revert to outer0). Next, we solve the outer problem (S68) with Dirichlet boundary conditions $\varrho(0)=\phi_{v},\varrho(L)=\phi_{l},\varrho_{a}(0)=\phi_{a,g}$ and $\varrho_{a}(L)=\phi_{a,l}$. In contrast to outer0, we now have $32(N-1)+$ 3 unknowns ($\varrho_{i},\varrho_{a,i}$ for $i=1,\dots,N-1$ and $c,A_{1},A_{2}$) and $12N+$ 1 equations [Eqs. (S70) for $i=0,\dots,N-1$ and one mass constraint (S71)]. The actives volume fraction $\phi_{a}$ is found a posteriori as $\phi_{a}=(1/N)\sum_{i=0}^{N-1}\varrho_{a,i}$.
@@
The nonlinear systems outer0,outer1,outer2 are again solved numerically using NonlinearSolve() with paramters reltol=1e-8, abstol=1e-8 and maxiters=20. We first initialise outer0 using the finite domain solution (Subsec. 2) with $\phi=0.80,\phi_{a}=0.45$, $\text{Pe}=7.5$, $L=100$, $N=3200$. Once an initial solution to outer0 has been found, we initialize the iterative solver using an existing solution with similar parameters: $\phi_{a}^{\prime}=\phi_{a}\pm 0.01$, $\phi_{p}^{\prime}=\phi_{p}\pm 0.01$. We first initialise outer1 using the finite domain solution (Subsec. 2) with $\phi=0.8,\phi_{a}=0.3$,$\text{Pe}=7.5$,$N=3200$. Once an initial solution to outer0 has been found, we initialize the iterative solver using an existing solution with similar parameters: $\phi^{\prime}=\phi\pm 0.01$, $\nu^{\prime}=\nu\pm 0.01$. We first initialise outer2 using the outer1 solution with $\phi=0.65,\nu=0.36$,$\text{Pe}=7.5$,$N=1024$, and place the new interface at $j=299$. Once an initial solution to outer2 has been found, we initialize the iterative solver using an existing solution with similar parameters: $\nu^{\prime}=\nu\pm 0.01$, $j^{\prime}=j\pm 8$.
Appendix F CP and T solutions

### v3

@@ 3 Traveling solutions in large systems
$\begin{aligned} -(c/L)\varrho+F&=A_{1},&\qquad F&=-\varrho^{\prime}-\tfrac{\text{Pe}^{2}}{2}(1-\varrho)\partial_{z}[d_{s}(\varrho)\varrho_{a}]\\ -(c/L)\varrho_{a}+F_{a}&=A_{2},&F_{a}&=-d_{s}(\varrho)\varrho_{a}^{\prime}-\varrho_{a}{\cal D}(\varrho)\varrho^{\prime}-\tfrac{\text{Pe}^{2}}{2}\left[\varrho_{a}s(\varrho)+d_{s}(\varrho)\right]\partial_{z}[d_{s}(\varrho)\varrho_{a}],\end{aligned}$ (S68)
@@
$\begin{aligned} -\frac{c}{2L}(\varrho_{i}+\varrho_{i+1})+F_{i+1/2}=A_{1},\qquad-\frac{c}{2L}(\varrho_{a,i}+\varrho_{a,i+1})+F_{a,i+1/2}=A_{2},\end{aligned}$ (S70)
@@
This yields a set of $2N+3$ unknowns ($\varrho_{i},\varrho_{a,i},c,A_{1},A_{2}$) and $2N+2$ equations [Eqs. (S70) and (S71)]. As in Subsec. 2, the additional constraint comes from noting that the problem has translational symmetry; we fix $\varrho_{0}=\phi$.
@@
Figure S5: Scattering in a counter-propagating solution. CP solution to (1). The initial condition is the homogeneous state ($\rho_{\pm}=\phi_{a}/2,\rho_{0}=\phi_{p}$) perturbed by a uniform random perturbation (see Sec. 1 for details). (a) Density profile at $t=4000$. (b) Kymographs showing the spatiotemporal dynamics. Parameters: $\text{Pe}=7.5$, $L=100.0$, $\phi_{a}=0.5$, $\phi_{p}=0.3$, $\Delta x=0.05$.
@@
Figure S6: Multi-stability of solutions. Two solutions, both obtained from Eq. (8) with the same parameters. (a) T solution. (b) PS solution. The top panels display the solution profile as a function of $z/L$. The bottom panels display the profile (solid black) overlayed on phase diagrams spanned by $\phi_{a}$ and $\phi_{p}$. The Spinodal (Blue) encloses the region of linear stability of homogeneous solutions, and the Binodal (Red) encloses the region of Phase separation. Black crosses mark the volume fraction. Parameters: $\text{Pe}=7.5$, $\phi_{a}=0.4608$, $\phi_{p}=0.2201$, $N=1024$
@@
Figure S7: Illustration of T and CP solutions. (a) Sketch of T solution with positive velocity, showing density of active and passive particles as a function of position (see Fig. 4 of main text for a similar solution with negative velocity). The red arrows indicate the self-propulsive force acting on the active particles. The blue arrows indicate forces applied by the active particles on the passive ones (via the cross-diffusion terms in the equation of motion). This generates the positive propagation. (b) Sketch of a CP solution at three different times in a large system (see Fig. 3 of main text for a similar solution in a smaller system). Initially the CP solution is well-approximated by a combination of two T solutions with opposite velocities of equal magnitude. The two propagating clusters approach each other and eventually collide. After the collision process one again sees a combination of two T solutions which travel away from each other. The cycle will eventually repeat due to periodic boundaries.
@@
For the outer1 problem with parameters $\phi$ and $\nu$, we first solve the coexisting phases problem (Sec. C) given $\nu$. This results in vapor and liquid phases $\phi_{v}(\nu)$ and $\phi_{l}(\nu)$, respectively, as well as individual components $(\varrho_{\sigma})_{v}$ and $(\varrho_{\sigma})_{l}$ in each phase. They satisfy $(\varrho_{\sigma})_{v}\equiv(\varrho_{+},\varrho_{-},\varrho_{0})_{v}=(\tfrac{1}{2}\phi_{a,g},\tfrac{1}{2}\phi_{a,g},\phi_{p,g})$ and similarly for $(\varrho_{\sigma})_{l}$. If $\phi_{v}(\nu)=\phi_{l}(\nu)$, it means there is no inner region and outer1 has no solution (and we revert to outer0). Next, we solve the outer problem (S68) with Dirichlet boundary conditions $\varrho(0)=\phi_{v},\varrho(L)=\phi_{l},\varrho_{a}(0)=\phi_{a,g}$ and $\varrho_{a}(L)=\phi_{a,l}$. In contrast to outer0, we now have 2(N1)+32(N-1)+3 unknowns ($\varrho_{i},\varrho_{a,i}$ for $i=1,\dots,N-1$ and $c,A_{1},A_{2}$) and $2N+1$ equations [Eqs. (S70) for $i=0,\dots,N-1$ and one mass constraint (S71)]. The actives volume fraction $\phi_{a}$ is found a posteriori as $\phi_{a}=(1/N)\sum_{i=0}^{N-1}\varrho_{a,i}$.
@@
The nonlinear systems outer0,outer1,outer2 are again solved numerically using NonlinearSolve() with paramters reltol=1e-8, abstol=1e-8 and maxiters=20. We first initialise outer0 using the finite domain solution (Subsec. 2) with $\phi=0.80,\phi_{a}=0.45$, $\text{Pe}=7.5$, $L=100$, $N=3200$. Once an initial solution to outer0 has been found, we initialize the iterative solver using an existing solution with similar parameters: $\phi_{a}^{\prime}=\phi_{a}\pm 0.01$, $\phi_{p}^{\prime}=\phi_{p}\pm 0.01$. We first initialise outer1 using the finite domain solution (Subsec. 2) with $\phi=0.8,\phi_{a}=0.3$$\text{Pe}=7.5$,$N=3200$. Once an initial solution to outer0 has been found, we initialize the iterative solver using an existing solution with similar parameters: $\phi^{\prime}=\phi\pm 0.01$, $\nu^{\prime}=\nu\pm 0.01$. We first initialise outer2 using the outer1 solution with $\phi=0.65,\nu=0.36$$\text{Pe}=7.5$,$N=1024$, and place the new interface at $j=299$. Once an initial solution to outer2 has been found, we initialize the iterative solver using an existing solution with similar parameters: $\nu^{\prime}=\nu\pm 0.01$, $j^{\prime}=j\pm 8$.
F CP and T solutions